A body is in simple harmonic motion with time period T=0.5s and amplitude A=1cm. Find the average velocity in the interval in which it moves from equilibrium position to half of its amplitude.
Ans 12cm/s
1
Expert's answer
2020-05-26T12:42:37-0400

Given that, a body in simple harmonic motion with time period "T=0.5s" and and amplitude is "A=1cm=10^{-2}m" .


The equation of simple harmonic motion will be,


"y=A\sin(\frac{2\pi}{T}t)\hspace{1cm}(\spades)"

Let at "t=t_1" ,the position of the body from equilibrium is "\frac{A}{2}" ,thus from above equation we get,


"\frac{A}{2}=A\sin(\frac{2\pi}{T}t_1)\\
\implies \sin(\frac{2\pi}{T}t_1)=\frac{1}{2}\\
\implies \frac{2\pi}{T}t_1=\frac{\pi}{6}\\
\implies t_1=\frac{T}{12}\hspace{1cm}(\clubs)"

Clearly, if we differentiate "(\spades)" w.r.t time we get the expression for velocity of the body,thus


"v=\frac{2\pi A}{T}\cos(\frac{2\pi }{T}t)"

Hence,

"<v>=\frac{1}{t_1-0}\int_{0}^{t_1}vdt\\
\implies <v>=\frac{1}{t_1-0}\int_{0}^{t_1}\frac{2\pi A}{T}\cos(\frac{2\pi }{T}t)dt\\
\implies <v>=\frac{2\pi A}{Tt_1}\int_{0}^{t_1}\cos(\frac{2\pi }{T}t)dt\\
\implies <v>=\frac{2\pi A}{Tt_1}\frac{T}{2\pi}\bigg[\sin(\frac{2\pi }{T}t)\bigg]_{0}^{t_1}\\
\implies <v>=\frac{A}{t_1}\sin(\frac{\pi }{6})=\frac{A}{2t_1}=12cm/s"